• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 14:06:14 on June 26, 2005

    Entry number 147260

    keyword=target update / shift schedule

    Dave Meekins and Gary Dezern are changing the target windows now. They have the chamber open and are mopping up the water (less than 1/4 cup, is the estimated volume). They expect to be done in a few hours with changing the window but the pumpdown time of the chamber+spectrometers is long (overnight).

    Day and swing shift workers are on standby, meaning you should leave a phone number with the counting house but only 1 person on the shift needs to be here. I will update these instructions after I hear from the technicians what the estimated time will be until we can have beam back.

    Most probable time for beam to return is early tomorrow morning.

    A copy of this log entry has been emailed to: lerose@jlab.org